### v1

@@ 1 Introduction
QMC methods are a class of quadrature rules designed to compute high-dimensional integrals (or equivalently expectations). The key benefits of QMC methods are that they can be tailored to achieve dimension-independent error bounds with faster than MC convergence rates for smooth integrands. QMC methods have demonstrated impressive performance in several UQ applications involving PDEs with random coefficients or domains, such as linear elliptic PDEs [12, 15, 18, 19, 20, 22, 29], elliptic eigenvalue problems [17, 37] and more recently linear parabolic equations with control [21]. However, much of the theory has focused on linear PDEs. The present work extends the theory in the case of non-linear, specifically semi-linear, parabolic equations, providing a roadmap for adapting QMC methods to wider classes of applications. Apart from the above work on eigenvalue problems [^3] While eigenvalue problems are also nonlinear, the differential operators [17, 37] are linear., to the best of our knowledge this paper is the first to analyze QMC methods for a non-linear PDE.
@@
However, the nonlinear term considered here, namely, u(1u)u(1-u) (see (2.1)), does not fit the settings in [7, 23, 24]. In particular, it is not monotone and so requires a novel proof, which uses the property that if our initial condition lies in $[0,1]$ then the solution remains in $[0,1]$ (proven in Section 2.5). We also makes use of the falling factorial combinatorial bound from [6]. We stress that obtaining regularity bounds for parabolic (time-dependent) nonlinear problems here is significantly harder than for elliptic problems in [6, 7, 23, 24] due to the need to handle both space and time derivatives and integrals in the norms. The parametric regularity bounds presented here are also of independent interest beyond QMC since they are often required to analyze other parametric approximation methods, e.g., sparse grids or stochastic collocation.
@@ 2.1 Governing equation
$f_{\rm rt}({\boldsymbol{x}},t)$ $=\gamma_{\text{rt},\varepsilon}\sum_{k}\big[1-\exp(-\alpha_{\text{rt}}\,z_{\text{rt},k}({\boldsymbol{x}})-\beta_{\text{rt}}\,z^{2}_{\text{rt},k}({\boldsymbol{x}}))\big]\,\,\mathbbold{1}(t\in[t_{\text{rt},k},t_{\text{rt},k}+\varepsilon]),$ (2.3)
@@
$\begin{cases}0<a_{\min}^{\boldsymbol{y}}\leq a^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq a_{\max}^{\boldsymbol{y}}<\infty,\\ 0\leq\kappa_{\min}^{\boldsymbol{y}}\leq\kappa^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq\kappa_{\max}^{\boldsymbol{y}}<\infty,\\ 0=f_{\min}\leq f({\boldsymbol{x}},t)\leq f_{\max}<\infty,\end{cases}\qquad\mbox{for all }{\boldsymbol{x}}\in\Omega,\;t\in I,\;{\boldsymbol{y}}\in U.$ (2.5)
@@ 2.2 Functional framework
$\|z\|_{H^{1}(\Omega)^{*}}:=\sup_{0\neq v\in H^{1}(\Omega)}\frac{|\langle z,v\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}|}{\|v\|_{H^{1}(\Omega)}}=\sup_{0\neq v\in H^{1}(\Omega)}\frac{|\int_{\Omega}z\,v\,{\mathrm{d}}{\boldsymbol{x}}|}{\|v\|_{H^{1}(\Omega)}}.$
@@
We follow Tröltzsch [45, §3.4] to analyze our PDE solution in the space $\mathcal{X}$} (denoted by $W(0,T)$ in [45]) which consists of all functions $v\in L^{2}(I;H^{1}(\Omega))$ with (distributional) derivative $\tfrac{\partial}{\partial t}v\in L^{2}(I;H^{1}(\Omega)^{*})$, and we equip the space $\mathcal{X}$} with the (graph) norm
@@
A fundamental property we will use is that the space $\mathcal{X}$} is embedded in $C(I;L^{2}(\Omega))$, see [45, Theorems 3.10 and 3.11], and therefore we formally write for $v\in{\mathcal{X}}$,
$\int_{I}\int_{\Omega}\frac{\partial v}{\partial t}v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=\int_{I}\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|v(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t=\frac{\|v(\cdot,T)\|_{L^{2}(\Omega)}^{2}-\|v(\cdot,0)\|_{L^{2}(\Omega)}^{2}}{2}.$
We will make use of a number of embedding constants which we summarize together here:
@@
$\|v\|_{C(I;L^{2}(\Omega))}$ $\leq\theta_{2}\,\|v\|_{\mathcal{X}}.$ (2.7)
The embedding (2.6) is due to the Sobolev embedding theorem for $H^{1}(\Omega)$, see e.g., [45, Theorem 7.1], which state that $H^{1}(\Omega)$ is embedded in $L^{p}(\Omega)$ when $d=2$ and $1\leq p<\infty$, and when $d\geq 3$ and $1\leq p\leq\frac{2d}{d-2}$.
Another property we will use is that the space $H^{1}(\Omega)$ is closed under addition, subtraction, and the maximum and minimum operations, see e.g., [16, Lemma 7.6]. This allows us to define continuous piecewise-smooth test functions in $H^{1}(\Omega)$.
@@ 2.3 Reparameterization
In the case where $f=0$, that is, untreated growth, the PDE model (2.1) is the well-studied Fisher-KPP equation [14, 28], and well-posedness results are readily available, see e.g., [39, §3.10.3]. However, we consider $f\neq 0$, thus standard results are not immediately applicable. PDE I (2.1) can be rewritten as
@@
Following a standard technique for treating parabolic PDEs involving a linear term with a negative coefficient (see e.g., [13, §7.5 Problem 8]), we use the substitution
@@
$\begin{cases}0<a_{\min}^{\boldsymbol{y}}\leq a^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq a_{\max}^{\boldsymbol{y}}<\infty,\\ 0\leq\kappa_{\min}^{\boldsymbol{y}}\leq b^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq\kappa_{\max}^{\boldsymbol{y}}\,e^{\lambda^{\boldsymbol{y}}\,T}<\infty,\\ 0<\lambda^{\boldsymbol{y}}-\kappa_{\max}^{\boldsymbol{y}}\leq c^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq\lambda^{\boldsymbol{y}}+f_{\max}-\kappa_{\min}^{\boldsymbol{y}}<\infty,\end{cases}\mbox{for all }{\boldsymbol{x}}\in\Omega,\;t\in I,\;{\boldsymbol{y}}\in U.$ (2.10)
@@ 2.4 Well-posedness
$\begin{cases}\displaystyle\int_{I}\int_{\Omega}\frac{\partial u}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=-\int_{I}\int_{\Omega}a\,\nabla u\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t-\int_{I}\int_{\Omega}(\kappa\,u^{2}+(f-\kappa)\,u)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t&\forall\,v\in{\mathcal{X}},\vskip 2.84544pt\\ \displaystyle\int_{\Omega}u(\cdot,0)\,v\,{\mathrm{d}}{\boldsymbol{x}}=\int_{\Omega}u_{0}\,v\,{\mathrm{d}}{\boldsymbol{x}}&\forall\,v\in L^{2}(\Omega).\end{cases}$ (2.11)
The first condition was obtained by multiplying the PDE with a test function $v\in{\mathcal{X}}$, integrating over ${\boldsymbol{x}}\in\Omega$ and $t\in I$, applying integration by parts on the diffusion term, and noting that $\int_{\partial\Omega}a\,\nabla u\cdot{\boldsymbol{n}}\,v\,{\mathrm{d}}{\boldsymbol{x}}=0$ due to the Neumann boundary condition.
@@
$\begin{cases}\displaystyle\int_{I}\int_{\Omega}\frac{\partial w}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=-\int_{I}\int_{\Omega}a\,\nabla w\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t-\int_{I}\int_{\Omega}(b\,w^{2}+c\,w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t&\forall\,v\in{\mathcal{X}},\vskip 2.84544pt\\ \displaystyle\int_{\Omega}w(\cdot,0)\,v\,{\mathrm{d}}{\boldsymbol{x}}=\int_{\Omega}w_{0}\,v\,{\mathrm{d}}{\boldsymbol{x}}&\forall\,v\in L^{2}(\Omega).\end{cases}$ (2.12)
@@
For every ${\boldsymbol{y}}\in U$ and $\lambda^{\boldsymbol{y}}>\kappa_{\max}^{\boldsymbol{y}}\geq 0$, consider PDE II (2.8) with bounded initial condition $w_{0}\in L^{2}(\Omega)$, $0\leq w_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$, and bounded coefficients (2.9)–(2.10). Then, there exists a unique weak solution $w^{\boldsymbol{y}}\in{\mathcal{X}}$ to PDE II, which is bounded by $0\leq w^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq e^{-\lambda^{\boldsymbol{y}}\,t}$ for a.e. ${\boldsymbol{x}}\in\Omega$ and all $t\in I$, and it satisfies the a priori bound $\|w^{\boldsymbol{y}}\|_{{\mathcal{X}}}\leq C^{\boldsymbol{y}}\,\|w_{0}\|_{L^{2}(\Omega)}$ with a constant $C^{\boldsymbol{y}}>0$.
@@
    For fixed ${\boldsymbol{y}}\in U$ and $\lambda^{\boldsymbol{y}}>\kappa_{\max}^{\boldsymbol{y}}$, we infer from [45, Assumptions 5.1 and 5.2, Lemmas 5.3 and 7.10] that there exists a unique weak solution $w_{*}^{\boldsymbol{y}}\in{\mathcal{X}}$ of PDE III with bounded coefficients (2.10), with a priori bound $\|w_{*}^{\boldsymbol{y}}\|_{{\mathcal{X}}}\leq C^{\boldsymbol{y}}\,\|w_{0}\|_{L^{2}(\Omega)}$ for $C^{\boldsymbol{y}}>0$.
@@
    Finally we prove that this solution $w_{*}^{\boldsymbol{y}}\in{\mathcal{X}}$ for PDE II is unique, and hence conclude that $u^{\boldsymbol{y}}({\boldsymbol{x}},t)=e^{\lambda^{\boldsymbol{y}}\,t}\,w_{*}^{\boldsymbol{y}}({\boldsymbol{x}},t)$ is the unique solution of PDE I, and it satisfies $0\leq u({\boldsymbol{x}},t)\leq 1$ for a.e. ${\boldsymbol{x}}\in\Omega$ and all $t\in I$.
The strategy of working with a cutoff function in PDE III and using piecewise test functions (see our proof below) appears already in the literature, sometimes referred to as Stampacchia truncation, see e.g., [4, 10, 39, 45]. However, the problem settings in those references differ from ours, so we include a proof for completeness.
@@
We see that PDE III satisfies [45, Assumption 5.1]: $\Omega$ is a bounded Lipschitz domain; for fixed $w\in{\mathbb{R}}$, the function $d({\boldsymbol{x}},t,w)$ is measurable with respect to ${\boldsymbol{x}}\in\Omega$ and $t\in I$; for all ${\boldsymbol{x}}\in\Omega$ and $t\in I$, the function $d({\boldsymbol{x}},t,w)$ is monotonically increasing (nondecreasing) with respect to $w\in{\mathbb{R}}$. We see that PDE III also satisfies [45, Assumption 5.2]: for all ${\boldsymbol{x}}\in\Omega$ and $t\in I$, the function $d({\boldsymbol{x}},t,w)$ is uniformly bounded and Lipschitz in $w\in{\mathbb{R}}$.
Hence, by [45, Lemma 5.3] and subsequent discussions, there exists a unique weak solution of PDE III in $\mathcal{X}$}. Consequently, by [45, Lemma 7.10], this solution satisfies $\|w^{\boldsymbol{y}}\|_{{\mathcal{X}}}\leq C^{\boldsymbol{y}}\,\|w_{0}\|_{L^{2}(\Omega)}$ with a constant $C^{\boldsymbol{y}}>0$.
We proceed to prove that this unique weak solution for PDE III is bounded between 0 and $e^{-\lambda\,t}$ if the initial condition is bounded by $0\leq w_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$. Integrating (2.13), using integration by parts on the diffusion term and using the zero Neumann boundary condition, we have for a.e. $t\in I$ and all $v\in H^{1}(\Omega)$,
$\int_{\Omega}\frac{\partial w}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}$ $=-\int_{\Omega}a\,\nabla w\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}d(w)\,v\,{\mathrm{d}}{\boldsymbol{x}}.$ (2.15)
Lower bound. We define $v=w_{\#}(\cdot,t)\in H^{1}(\Omega)$ to be the “unwanted lower part” of $w$,
@@
$\frac{\partial w}{\partial t}\,w_{\#}=\frac{\partial w_{\#}}{\partial t}\,w_{\#},\quad\nabla w\cdot\nabla w_{\#}=|\nabla w_{\#}|^{2},\quad d(w)\,w_{\#}=0,$ (2.16)
where we used Case 1 of (2.14). On the other hand, if $w({\boldsymbol{x}},t)\geq 0$ then $w_{\#}({\boldsymbol{x}},t)=0$ and the equalities in (2.16) hold trivially. Putting (2.16) into (2.15) gives for a.e. $t\in I$,
$\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}^{2}=\int_{\Omega}\frac{\partial w_{\#}}{\partial t}\,w_{\#}\,{\mathrm{d}}{\boldsymbol{x}}=-\int_{\Omega}a\,|\nabla w_{\#}|^{2}\,{\mathrm{d}}{\boldsymbol{x}}\leq 0.$
We conclude that $\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}$ is a nonincreasing function of $t$. Thus $\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}\leq\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}=0$, and hence $w_{\#}({\boldsymbol{x}},t)=0$ for a.e. ${\boldsymbol{x}}\in\Omega$, and in turn this yields
$w({\boldsymbol{x}},t)\geq 0\qquad\mbox{for a.e.\ }{\boldsymbol{x}}\in\Omega\mbox{ and all }t\in I.$
Upper bound. Now we define $v=w_{\#}(\cdot,t)\in H^{1}(\Omega)$ on the “unwanted upper part” of $w$,

### v2

@@ 1 Introduction
QMC methods are a class of quadrature rules designed to compute high-dimensional integrals (or equivalently expectations). The key benefits of QMC methods are that they can be tailored to achieve dimension-independent error bounds with faster than MC convergence rates for smooth integrands. QMC methods have demonstrated impressive performance in several UQ applications involving PDEs with random coefficients or domains, such as linear elliptic PDEs [13, 17, 20, 21, 22, 24, 32], elliptic eigenvalue problems [19, 40] and more recently linear parabolic equations with control [23]. However, much of the theory has focused on linear PDEs. The present work extends the theory in the case of non-linear, specifically semi-linear, parabolic equations, providing a roadmap for adapting QMC methods to wider classes of applications. To the best of our knowledge, this paper is the first to analyze QMC methods for a non-linear PDE.
@@
However, the nonlinear term considered here, namely, u(1u)u(1-u) (see (2.1)), does not fit the settings in [7, 25, 26]. In particular, it is not monotone and so requires a novel proof, which uses the property that if our initial condition lies in $[0,1]$ then the solution remains in $[0,1]$ (proven in Section 2.5). We also make use of the falling factorial combinatorial bound from [6]. We stress that obtaining regularity bounds for parabolic (time-dependent) nonlinear problems here is significantly harder than for elliptic problems in [6, 7, 25, 26] due to the need to handle both space and time derivatives and integrals in the norms. The parametric regularity bounds presented here are also of independent interest beyond QMC since they are often required to analyze other parametric approximation methods, e.g., sparse grids or stochastic collocation.
@@ 2.1 Governing equation
$f_{\rm rt}({\boldsymbol{x}},t)$ $=\gamma_{\text{rt},\varepsilon}\sum_{k}\big[1-\exp(-\alpha_{\text{rt}}\,z_{\text{rt},k}({\boldsymbol{x}})-\beta_{\text{rt}}\,z^{2}_{\text{rt},k}({\boldsymbol{x}}))\big]\,\,\mathbbold{1}(t\in[\tau_{\text{rt},k},\tau_{\text{rt},k}+\varepsilon]),$ (2.3)
@@
$\begin{cases}0<a_{\min}^{\boldsymbol{y}}\leq a^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq a_{\max}^{\boldsymbol{y}}<\infty,\\ 0\leq\,\kappa_{\min}^{\boldsymbol{y}}\leq\kappa^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq\kappa_{\max}^{\boldsymbol{y}}<\infty,\;\kappa_{\max}^{\boldsymbol{y}}>0,\\ 0=f_{\min}\leq\,f({\boldsymbol{x}},t)\leq f_{\max}<\infty,\end{cases}\qquad\mbox{for all }{\boldsymbol{x}}\in\Omega,\;t\in I,\;{\boldsymbol{y}}\in U.$ (2.5)
@@ 2.2 Functional framework
$\|z\|_{H^{1}(\Omega)^{*}}:=\sup_{0\neq v\in H^{1}(\Omega)}\frac{|\langle z,v\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}|}{\|v\|_{H^{1}(\Omega)}}.$
@@
We follow Tröltzsch [49, §3.4] to analyze our PDE solution in the space $\mathcal{X}$} (denoted by $W(0,T)$ in [49]) which consists of all functions $v\in L^{2}(I;H^{1}(\Omega))$ with (distributional) derivative $\tfrac{\partial}{\partial t}v\in L^{2}(I;H^{1}(\Omega)^{*})$, and we equip the space $\mathcal{X}$} with the (graph) norm
@@
A fundamental property we will use is that the space $\mathcal{X}$} is embedded in $C(I;L^{2}(\Omega))$, see [49, Theorems 3.10 and 3.11]. We formally write, for $v\in{\mathcal{X}}$ and all $\tau\in I$
$\int_{0}^{\tau}\!\big\langle\tfrac{\partial}{\partial t}v,v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t=\int_{0}^{\tau}\!\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|v(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t=\frac{\|v(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}-\|v(\cdot,0)\|_{L^{2}(\Omega)}^{2}}{2}.$ (2.6)
We will make use of a number of embeddings which we summarize together here:
@@
$\|v\|_{C(I;L^{2}(\Omega))}$ $\leq\theta_{2}\,\|v\|_{\mathcal{X}},$ (2.8)
with $\theta_{1},\theta_{2}>0$. The embedding (2.7) is due to the Sobolev embedding theorem for $H^{1}(\Omega)$, see e.g., [49, Theorem 7.1], which states that $H^{1}(\Omega)$ is embedded in $L^{p}(\Omega)$ when $d=2$ and $1\leq p<\infty$, and when $d\geq 3$ and $1\leq p\leq\frac{2d}{d-2}$.
Another property we will use is that the space $H^{1}(\Omega)$ is closed under addition, subtraction, and the maximum and minimum operations, see e.g., [18, Lemma 7.6]. This allows us to define continuous piecewise-smooth test functions in $H^{1}(\Omega)$.
@@ 2.3 Reparameterization
In the case where $f=0$, that is, untreated growth, the PDE model (2.1) is the well-studied Fisher-KPP equation [15, 30], and well-posedness results are readily available, see e.g., [42, §3.10.3]. However, we consider $f\neq 0$, thus standard results are not immediately applicable. PDE I (2.1) can be rewritten as
@@
Following a standard technique for treating parabolic PDEs involving a linear term with a negative coefficient (see e.g., [14, §7.5 Problem 8]), we use the substitution
@@
$\begin{cases}0<a_{\min}^{\boldsymbol{y}}\leq a^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq a_{\max}^{\boldsymbol{y}}<\infty,\\ 0\leq\kappa_{\min}^{\boldsymbol{y}}\leq b_{\min}^{\boldsymbol{y}}\leq b^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq b_{\max}^{\boldsymbol{y}}\leq\kappa_{\max}^{\boldsymbol{y}}\,e^{\lambda^{\boldsymbol{y}}\,T}<\infty,\\ 0<\lambda^{\boldsymbol{y}}-\kappa_{\max}^{\boldsymbol{y}}\leq c_{\min}^{\boldsymbol{y}}\leq c^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq c_{\max}^{\boldsymbol{y}}\leq\lambda^{\boldsymbol{y}}+f_{\max}<\infty,\end{cases}\mbox{for all }{\boldsymbol{x}}\in\Omega,\;t\in I,\;{\boldsymbol{y}}\in U.$ (2.11)
@@ 2.4 Well-posedness
$\begin{cases}\displaystyle\int_{I}\big\langle\tfrac{\partial}{\partial t}u,v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t+\!\int_{I}\!\int_{\Omega}\!\big(a\,\nabla u\cdot\nabla v+(\kappa\,u^{2}+(f-\kappa)\,u)\,v\big)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=0&\forall\,v\in{\mathcal{X}},\vskip 2.84544pt\\ \displaystyle\int_{\Omega}u(\cdot,0)\,v\,{\mathrm{d}}{\boldsymbol{x}}=\int_{\Omega}u_{0}\,v\,{\mathrm{d}}{\boldsymbol{x}}&\forall\,v\in L^{2}(\Omega).\end{cases}$ (2.12)
The first condition was obtained by testing the PDE with a function $v\in{\mathcal{X}}$, integrating over ${\boldsymbol{x}}\in\Omega$ and $t\in I$, applying integration by parts on the diffusion term, and noting that $\int_{\partial\Omega}a\,\nabla u\cdot{\boldsymbol{n}}\,v\,{\mathrm{d}}{\boldsymbol{x}}=0$ due to the Neumann boundary condition. The integral $\int_{I}\int_{\Omega}u^{2}\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$ is well defined because if $u\in{\mathcal{X}}$ then $u^{2}\in L^{2}(I;L^{2}(\Omega))$ via the embedding (2.7).
@@
$\begin{cases}\displaystyle\int_{I}\big\langle\tfrac{\partial}{\partial t}w,v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}\big(a\,\nabla w\cdot\nabla v+(b\,w^{2}+c\,w)\,v\big)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=0&\forall\,v\in{\mathcal{X}},\vskip 2.84544pt\\ \displaystyle\int_{\Omega}w(\cdot,0)\,v\,{\mathrm{d}}{\boldsymbol{x}}=\int_{\Omega}w_{0}\,v\,{\mathrm{d}}{\boldsymbol{x}}&\forall\,v\in L^{2}(\Omega).\end{cases}$ (2.13)
@@
For every ${\boldsymbol{y}}\in U$ and $\lambda^{\boldsymbol{y}}>\kappa_{\max}^{\boldsymbol{y}}>0$, consider PDE II (2.9) with bounded initial condition $w_{0}\in L^{2}(\Omega)$, $0\leq w_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$, and bounded coefficients (2.10)–(2.11). Then, there exists a unique weak solution $w^{\boldsymbol{y}}\in{\mathcal{X}}$ to PDE II. This unique solution is bounded by $0\leq w^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq e^{-\lambda^{\boldsymbol{y}}\,t}$ for a.e. ${\boldsymbol{x}}\in\Omega$ and all $t\in I$, and it satisfies the a priori bound $\|w^{\boldsymbol{y}}\|_{{\mathcal{X}}}\leq C^{\boldsymbol{y}}\,\|w_{0}\|_{L^{2}(\Omega)}$ with a constant $C^{\boldsymbol{y}}>0$.
@@
    For fixed ${\boldsymbol{y}}\in U$ and $\lambda^{\boldsymbol{y}}>\kappa_{\max}^{\boldsymbol{y}}>0$, we infer from [49, Assumptions 5.1 and 5.2, Lemmas 5.3 and 7.10] that there exists a unique weak solution $w_{*}^{\boldsymbol{y}}\in{\mathcal{X}}$ of PDE III with bounded coefficients (2.11), with a priori bound $\|w_{*}^{\boldsymbol{y}}\|_{{\mathcal{X}}}\leq C^{\boldsymbol{y}}\,\|w_{0}\|_{L^{2}(\Omega)}$ for $C^{\boldsymbol{y}}>0$.
@@
    Finally we prove that this solution $w_{*}^{\boldsymbol{y}}\in{\mathcal{X}}$ for PDE II is unique. In particular, there is no other solution taking values below 0 or above $e^{-\lambda^{\boldsymbol{y}}\,t}$. Hence we conclude that $u^{\boldsymbol{y}}({\boldsymbol{x}},t)=e^{\lambda^{\boldsymbol{y}}\,t}\,w_{*}^{\boldsymbol{y}}({\boldsymbol{x}},t)$ is the unique solution of PDE I, and it satisfies $0\leq u({\boldsymbol{x}},t)\leq 1$ for a.e. ${\boldsymbol{x}}\in\Omega$ and all $t\in I$.
The strategy of working with a cutoff function in PDE III and using continuous piecewise-smooth test functions (see our proof below) appears already in the literature, sometimes referred to as Stampacchia truncation, see e.g., [4, 10, 42, 49]. However, the problem settings in those references differ from ours, so we include a proof for completeness.
@@
We see that PDE III satisfies [49, Assumption 5.1]: $\Omega$ is a bounded Lipschitz domain; for fixed $w\in{\mathbb{R}}$, the function $d({\boldsymbol{x}},t,w)$ is measurable with respect to ${\boldsymbol{x}}\in\Omega$ and $t\in I$; for all ${\boldsymbol{x}}\in\Omega$ and $t\in I$, the function $d({\boldsymbol{x}},t,w)$ is monotonically increasing (nondecreasing) with respect to $w\in{\mathbb{R}}$. We see that PDE III also satisfies [49, Assumption 5.2]: for all ${\boldsymbol{x}}\in\Omega$ and $t\in I$, the function $d({\boldsymbol{x}},t,w)$ is uniformly bounded and Lipschitz in $w\in{\mathbb{R}}$.
Hence, by [49, Lemma 5.3] and subsequent discussions, there exists a unique weak solution of PDE III in $\mathcal{X}$}. Consequently, by [49, Lemma 7.10], this solution satisfies $\|w^{\boldsymbol{y}}\|_{{\mathcal{X}}}\leq C^{\boldsymbol{y}}\,\|w_{0}\|_{L^{2}(\Omega)}$ with a constant $C^{\boldsymbol{y}}>0$.
We proceed to prove that this unique weak solution for PDE III is bounded between 0 and $e^{-\lambda\,t}$ if the initial condition is bounded by $0\leq w_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$. Integrating (2.14) with respect to ${\boldsymbol{x}}\in\Omega$, using integration by parts on the diffusion term and using the zero Neumann boundary condition, and then integrating over $t\in[0,\tau]$ for some $\tau\leq T$, we obtain
$\int_{0}^{\tau}\!\big\langle\tfrac{\partial}{\partial t}w,v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t+\int_{0}^{\tau}\!\int_{\Omega}\big(a\,\nabla w\cdot\nabla v+d(w)\,v\big)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=0\quad\forall\,v\in{\mathcal{X}}.$ (2.16)
Lower bound. We define $v=w_{\#}\in{\mathcal{X}}$ to be the “unwanted lower part” of $w$,
@@
$\tfrac{\partial}{\partial t}w=\tfrac{\partial}{\partial t}w_{\#},\quad\nabla w\cdot\nabla w_{\#}=|\nabla w_{\#}|^{2},\quad d(w)\,w_{\#}=0,$ (2.17)
where we used Case 1 of (2.15). On the other hand, if $w({\boldsymbol{x}},t)\geq 0$ then $w_{\#}({\boldsymbol{x}},t)=0$ and the equalities in (2.17) hold trivially. Putting (2.17) into (2.16) and using (2.6) gives for all $\tau\in I$,
$\frac{\|w_{\#}(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}-\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}^{2}}{2}$
$=\int_{0}^{\tau}\!\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t=\int_{0}^{\tau}\!\big\langle\tfrac{\partial}{\partial t}w_{\#},w_{\#}\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t=-\int_{0}^{\tau}\!\int_{\Omega}a\,|\nabla w_{\#}|^{2}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t\leq 0.$
Thus $\|w_{\#}(\cdot,\tau)\|_{L^{2}(\Omega)}\leq\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}=0$, and hence $w_{\#}({\boldsymbol{x}},\tau)=0$ for a.e. ${\boldsymbol{x}}\in\Omega$, and in turn this yields
$w({\boldsymbol{x}},\tau)\geq 0\qquad\mbox{for a.e.\ }{\boldsymbol{x}}\in\Omega\mbox{ and all }\tau\in I.$
Upper bound. Now we define $v=w_{\#}\in{\mathcal{X}}$ on the “unwanted upper part” of $w$,